# Efficient Image Pre-Training with Siamese Cropped Masked Autoencoders

### v1

@@ 4.4 Learning Speed
We evaluate the evolution of the performances of CropMAE and SiamMAE. In particular, we compare SiamMAE trained on K400, CropMAE trained on K400 and CropMAE trained on ImageNet Subset, all for $400$ epochs. The performance on the DAVIS-2017 object propagation task [39] is reported every $50$ epochs in Figure 6. Remarkably, our approach demonstrates superior performance when trained on the ImageNet Subset compared to training using K400 video frames. This improvement can be attributed to two main factors: (i) the greater diversity of the ImageNet dataset, containing a broader spectrum of objects, and (ii) its focus on currated object-centric images, which likely results in more relevant crops and reconstruction tasks. In contrast, random cropping in K400 frequently yields images without any objects, diminishing the effectiveness of the learning process.
@@
Figure 6: Performances of CropMAE and SiamMAE on DAVIS during pre-training. For a fixed number of $400$ epochs, CropMAE trains faster and yields better results than SiamMAE [23], when trained on K400 frames or ImageNet Subset images.
@@ 4.5 Training time
We compare the training times of CropMAE and SiamMAE. On the one hand, CropMAE uses an extremely high masking ratio, and only needs a single frame of a video clip to train, or even a standalone image. On the other hand, SiamMAE uses a masking ratio that is lower, and needs two different frames to work. Both these factors significantly impact the training time, as seeking distant frames may require decoding a larger portion of the video, and the number of operations performed by the attention layers increases quadratically with the number of visible patches [25]. We measure the total time taken by both approaches to train and report our results in Table 2. As can be seen, CropMAE trains almost $30\%$ faster than SiamMAE on K400 for a fixed computational budget, thanks to its use of fewer patches and frames. In addition, CropMAE trained on ImageNet, which does not require decoding videos, achieves a $2380\%$ speed-up on our hardware, while reaching better performances.
@@
Method Dataset Number of images Mask Ratio Speedup
SiamMAE K400 2 95% $\times$1.0
CropMAE K400 1 98.5% $\times$1.29
CropMAE IN Subset 1 98.5% $\times$23.8
@@ 4.6 Ablation Studies
We perform several ablation studies on the different components of CropMAE and report the results in Table 6(e). Unless stated otherwise, we use the default parameters (refer to the supplementary material). Specifically, we train CropMAE on our ImageNet subset for $400$ epochs and report the results obtained on the DAVIS-2017 [39] object propagation task.
@@
4.6.1 Cropping Strategy.
We study the effect on performance of different cropping strategies in Table LABEL:fig:crop_strategy_results. We can see that reconstructing the same views (Figure 2(a)) yields very poor performances ($36.6$), suggesting that the model failed to learn any propagation capabilities. Reconstructing the global view from the local view (Figure 2(c)) results in significantly improved performance ($55.9$) The Random Views (Figure 2(b)) and Global-to-Local (Figure 2(d)) approaches achieve the highest scores ($60.0$ and $60.4$, respectively). Interestingly, these setups are the only ones enabling a completely tractable task without any prior knowledge, meaning the reconstruction can solely rely on the unmasked image. In fact, tractability is sometimes guaranteed in the random setting, while it is always true for the Global-to-Local approach, which likely explains its superior performance.
4.6.2 Masking Ratio.
@@ 5 Conclusion
In this work, we introduce CropMAE, a self-supervised method for quickly learning rich features for video propagation tasks by reconstructing a crop of an image that has been masked at an extremely high proportion (over 98.5%). We empirically demonstrate that our method can learn useful features for video downstream tasks without requiring explicit video motion. These features can be learned from still images, resulting in even richer information. Thanks to our explicit pretext task, our method trains faster than existing methods and is applicable to both video frames and still images. Finally, we show on-par performances with state-of-the-art methods for three video propagation downstream tasks.
Limitations and future work. Despite being designed to work with small quantities of data and facilitate fast training, we believe the scalability of our method warrants further investigation. This includes both model scalability (i.e., patch size and ViT size) and data scalability (i.e., the amount of data available and the differences between images and video frames). More effort should be directed towards understanding the unique contributions of video frames instead of still images, especially concerning scalability, and determining their necessity to develop rich and robust representations.
@@ Acknowledgements
A. Cioppa is funded by the F.R.S.-FNRS. This work was supported by the KAUST OSR through the VCC funding and the SDAIA-KAUST Center of Excellence in Data Science and Artificial Intelligence. The present research benefited from computational resources made available on Lucia, the Tier-1 supercomputer of the Walloon Region, infrastructure funded by the Walloon Region under the grant agreement n°1910247. We acknowledge EuroCC Belgium for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium.

### v2

@@ 4.4 Learning Speed
We evaluate the evolution of the performances of CropMAE and SiamMAE. In particular, we compare SiamMAE trained on K400, CropMAE trained on K400, and CropMAE trained on ImageNet Subset, all for $400$ epochs. The performance on the DAVIS-2017 object propagation task [39] is reported every $50$ epochs in Figure 6. Remarkably, our approach demonstrates superior performance when trained on the ImageNet Subset compared to training using K400 video frames. This improvement can be attributed to two main factors: (i) the greater diversity of the ImageNet dataset, containing a broader spectrum of objects, and (ii) its focus on currated object-centric images, which likely results in more relevant crops and reconstruction tasks. In contrast, random cropping in K400 frequently yields images without any objects, diminishing the effectiveness of the learning process.
@@
Figure 6: Performances of CropMAE and SiamMAE on DAVIS during pre-training. For a fixed number of $400$ epochs, CropMAE trains faster and consistently yields better results than SiamMAE [23], when trained on K400 frames or ImageNet Subset images.
@@ 4.5 Training time
We compare the training times of CropMAE and SiamMAE. On the one hand, CropMAE uses an extremely high masking ratio, and only needs a single frame of a video clip to train, or even a standalone image. On the other hand, SiamMAE uses a lower masking ratio and needs two different frames to work. Both these factors significantly impact the training time, as seeking distant frames may require decoding a larger portion of the video, and the number of operations performed by the attention layers increases quadratically with the number of visible patches [25]. We measure the total time taken by both approaches to train and report our results in Table 2. As it can be seen, CropMAE trains almost $30\%$ faster than SiamMAE on K400 for a fixed computational budget, thanks to its use of fewer patches and frames. When pre-training on images (i.e., on the IN Subset), which are significantly faster to decode, CropMAE achieves a tremendous speed-up of 2380% on our hardware while also reaching better performances.
@@
Method Dataset Number of images Mask Ratio GFLOPS Speedup
SiamMAE K400 2 95% 5.8 $\times$1.0
CropMAE K400 1 98.5% 5.6 $\times$1.29
CropMAE IN Subset 1 98.5% 5.6 $\times$23.8
@@ 4.6 Ablation Studies
We perform several ablation studies on the different components of CropMAE and report the results in Table 6(e). Unless stated otherwise, we use the default parameters presented in the Appendix 6.1. Specifically, we train CropMAE on our ImageNet subset for $400$ epochs and report the results obtained on the DAVIS-2017 [39] object propagation task.
@@
Cropping Strategy.
We study the effect on performance of different cropping strategies in Table LABEL:fig:crop_strategy_results. We can see that reconstructing the same views (Figure 2(a)) yields very poor performances ($36.6$), suggesting that the model failed to learn any propagation capabilities. Reconstructing the Local-to-Global view (Figure 2(c)) results in significantly improved performance ($55.9$). The Random Views (Figure 2(b)) and Global-to-Local (Figure 2(d)) approaches achieve the highest scores ($60.0$ and $60.4$, respectively). Interestingly, these setups are the only ones enabling a completely tractable task without any prior knowledge, meaning the reconstruction can solely rely on the unmasked image. In fact, tractability is sometimes guaranteed in the random setting, while it is always true for the Global-to-Local approach, which likely explains its superior performance.
Masking Ratio.
@@ 5 Conclusion
In this work, we introduce CropMAE, a self-supervised method for quickly learning rich features for video propagation tasks by reconstructing a crop of an image that has been masked at an extremely high proportion (over 98.5%). We empirically demonstrate that our method can learn useful features for video downstream tasks without requiring explicit video motion. These features can be learned from still images, resulting in even richer information. Thanks to our tractable pretext task, our method trains faster than existing methods and is applicable to both video frames and still images. Finally, we show on-par performances with state-of-the-art methods for three video propagation downstream tasks.
Limitations and future work.
Despite being designed to work with small quantities of data and facilitate fast training, we believe the scalability of our method warrants further investigation. This includes both model scalability (i.e., patch size and ViT size) and data scalability (i.e., the amount of data available and the differences between images and video frames). More effort should be directed towards understanding the unique contributions of video frames instead of still images, especially concerning scalability, and determining their necessity to develop rich and robust representations.
@@ Acknowledgements
A. Cioppa is funded by the F.R.S.-FNRS. The research reported in this publication was supported by funding from KAUST Center of Excellence on GenAI, under award number 5940, and the SDAIA-KAUST Center of Excellence in Data Science and Artificial Intelligence. The present research benefited from computational resources made available on Lucia, the Tier-1 supercomputer of the Walloon Region, infrastructure funded by the Walloon Region under the grant agreement n°1910247. We acknowledge EuroCC Belgium for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium.